# Wenjun Huang1  Jianguo Hu1

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have attracted much attention due to their multifunctionality. However, traditional Transformer architectures incur significant overhead due to their secondary computational complexity. To address this issue, we introduce ML-Mamba, a multimodal language model that utilizes the latest and efficient Mamba-2 model for inference. Mamba-2 is known for its linear extension and fast processing of long sequences. We replace the Transformer based backbone with a pre-trained Mamba-2 model and explore methods for integrating 2D visual selective scanning mechanisms into multimodal learning. We also try various visual encoders and Mamba-2 model variants. Our extensive experiments conducted in various multimodal benchmark tests have demonstrated the competitive performance of ML-Mamba and highlighted the potential of state space models in multimodal tasks. The experimental results show that: (1) ML-Mamba achieves performance comparable to state-of-the-art methods such as TinyLaVA and MobileVLM v2 through its linear sequential modeling, while also having faster inference speed; (2) ML-Mamba performs well in visual hallucinations and spatial relationship judgment in closed set benchmark tests; (3) ML-Mamba achieves performance comparable to LLaVA while reducing the number of parameters by 40%.(4) Compared to the multimodal model using the original Mamba model, the Mamba-2 based large-scale multimodal language model has stronger inference performance and effectiveness.
@@ 1 Introduction
Figure 1: Illustration of tokens per second and times in our proposed ML-Mamba and baselines.
@@
To this end, this article proposes a new perspective, which is to directly use the state space model (SSM) as the backbone. Specifically, we use the Mamba-2 language model as the basic model of our VLM. In this article, we introduce ML-Mamba, a work that applies state space models to multimodal learning tasks. Our method utilizes a pre-trained Mamba-2 language model as the backbone, replacing traditional Transformer based models such as LLaMA [50] or Vicuna [4]. We further enhanced ML-Mamba through a novel multimodal connector called Mamba-2 Scan Connector (MSC) architecture, which includes a Mamba-2 visual selective scanning module (MVSS) and a SwiGLU module specifically designed for 2D causal modeling of enriched visual sequences. The MVSS module explores two different scanning mechanisms: bidirectional scanning mechanism (BSM) and cross scanning mechanism (CSM). In addition, we investigated the combination of different visual encoders, variants of pre-trained Mamba-2 language models, and multimodal connectors to optimize the integration of visual and linguistic information.
Extensive experiments conducted on various multimodal learning benchmarks have demonstrated the effectiveness of ML-Mamba. Our model achieves competitive performance with other similarly sized small MLLMs, and even outperforms large MLLMs on some popular benchmark tests (such as LLaVA v1.5 [34] versions 7B and 13b).
In summary, our contributions are as follows:
  • We propose ML-Mamba, which is a work that explores and utilizes multimodal learning tasks combined with the latest Mamba-2. Compared with the multimodal model using the original Mamba, the multimodal large-scale language model based on Mamba-2 has stronger inference performance and effectiveness. ML-Mamba also provides a new framework choice for multimodal large-scale language models beyond Transformer based architectures.
  • We have empirically explored the impact of different components in ML-Mamba and proposed a novel multi-mode connector called Mamba-2 Scan Connector (MSC), which includes Mamba-2 Visual Selective Scanning (MVSS) module and SwiGLU module to enhance characterization capability.
  • We conducted extensive experiments on different multimodal learning benchmarks and demonstrated that ML-Mamba achieves competitive performance compared to existing multimodal large-scale language models.
  • We will open source the code to promote research on applying state space models to multimodal learning.
@@ 3.1 Mamba Preliminaries
By combining the modified parallel Mamba blocks with using SSD as the inner SSM layer, the Mamba-2 architecture is formed (as shown on the left in Fig. 2). The performance of Mamba-2 models of varying sizes on the Pile dataset shows that it matches or outperforms Mamba and other open-source Transformer models on standard downstream evaluations.
Figure 2: The architecture of ML-Mamba (right) uses Mamba-2 as the backbone (left). It includes a visual encoder, a multi-modal connector called the Mamba-2 Scan Connector (MSC), a MLP projector, and a language model. We use the pre-trained Mamba-2 large language model (Mamba-2 LLM) as the language model and a pre-trained visual transformer model as the visual encoder.
@@ 3.2.3 MultiModal Connector
  • MSC-MLP (Basic): It combines the multimodal connector called the Mamba-2 Scan Connector (MSC) module, which does not include the SwiGLU module and is intended to enhance the processing capability of two-dimensional non-causal visual information. Subsequently, the MLP aligns the features of vision and text (see Fig. 3(b)).
@@ 3.2.5 Training Process
The recent research [24] indicates that for existing training paradigms based on LLaVA [6, 34, 57] (i.e., training only the projection layer in the pre-alignment stage and fine-tuning the LLM backbone in one iteration each), we decided to first use a 558K subset of the LAION-CC-SBU dataset to align the Mamba-2 Scan Connector (MSC) and the projector. During the fine-tuning stage, we simultaneously optimized the Mamba-2 Scan Connector (MSC), the projector, and the Mamba LLM. This comprehensive training effort was executed on 8 NVIDIA Tesla A100 GPUs. We conducted fine-tuning over two randomly sampled epochs using a composite dataset, which includes:
1. LVIS-Instruct-4V [51]: This dataset contains 220K images with visual alignment and context-aware instructions generated by GPT-4V.
2. The Mixed Dataset Used in LLaVA v1.5: This dataset includes a total of 655K visual multi-round dialogue samples. It encompasses academic VQA samples [13, 21, 26, 46], as well as visual instruction tuning data from LLaVA Instruction [45] and ShareGPT [35].
Overall, the fine-tuning dataset consists of approximately 875K images and their corresponding multi-turn dialogue data, as well as pure text dialogue data.
@@ 4.1 Experimental Setup
Projector init Random
Mamba-2 Scan Connector init Random
@@
number of samples in the instruction tuning stage 558K
number of samples in the pretraining stage 875K
@@
Number of epochs 2
@@
ML-Mamba (ours) Mamba-2 LLM-2.7B 558K 875K 75.26 60.68 52.2 88.3 45.17 51.5
@@ 4.2 Results
To demonstrate the effectiveness of the model, we compared it with a VLM of the same scale with approximately 3B parameters, or with a larger VLM containing twice the number of parameters. As shown in Table  2, although Although ML-Mamba has only about 40% of its parameters, it performs well on multiple benchmarks compared to LLaVA v1.5 7B, and even outperforms all models on POPE.
@@ 4.3 Reasoning speed
The results from Fig. 1 and Table 3 demonstrated that although the number of image markers processed by ML-Mamba significantly increased, it still exhibited extremely fast inference speed. Compared to MobileVLM v2, although the latter has undergone multiple lightweight optimizations, the time required for ML-Mamba to complete inference is only about 30% of the former. This indicates that ML-Mamba not only maintains high speed while processing larger data, but also, thanks to the characteristics of its RNN like model, its memory usage does not significantly increase with the increase of image marker length, as such models maintain a fixed size hidden state to store historical information during the inference process.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have attracted much attention for their multifunctionality. However, traditional Transformer architectures incur significant overhead due to their secondary computational complexity. To address this issue, we introduce ML-Mamba, a multimodal language model, which utilizes the latest and efficient Mamba-2 model for inference. Mamba-2 is known for its linear scalability and fast processing of long sequences. We replace the Transformer-based backbone with a pre-trained Mamba-2 model and explore methods for integrating 2D visual selective scanning mechanisms into multimodal learning while also trying various visual encoders and Mamba-2 model variants. Our extensive experiments in various multimodal benchmark tests demonstrate the competitive performance of ML-Mamba and highlight the potential of state space models in multimodal tasks. The experimental results show that: (1) we empirically explore how to effectively apply the 2D vision selective scan mechanism for multimodal learning. We propose a novel multimodal connector called the Mamba-2 Scan Connector (MSC), which enhances representational capabilities. (2) ML-Mamba achieves performance comparable to state-of-the-art methods such as TinyLaVA and MobileVLM v2 through its linear sequential modeling while faster inference speed; (3) Compared to multimodal models utilizing Mamba-1, the Mamba-2-based ML-Mamba exhibits superior inference performance and effectiveness.
@@ 1 Introduction
To this end, this article proposes a new perspective, directly using the state space model (SSM) as the backbone. Specifically, we use the Mamba-2 language model as the basic model of our VLM. In this article, we introduce ML-Mamba, a work that applies state space models to multimodal learning tasks. Our method utilizes a pre-trained Mamba-2 language model as the backbone, replacing traditional Transformer-based models such as LLaMA [46]. We further enhanced ML-Mamba through a novel multimodal connector called Mamba-2 Scan Connector (MSC) architecture, which includes a Mamba-2 visual selective scanning module (MVSS) and a SwiGLU module specifically both designed for 2D causal modeling of enriched visual sequences. The MVSS module explores two different scanning mechanisms: bidirectional scanning mechanism (BSM) and cross scanning mechanism (CSM). In addition, we investigated the combination of different visual encoders, variants of pre-trained Mamba-2 language models, and multimodal connectors to optimize the integration of visual and linguistic information.
Extensive experiments conducted on a range of multimodal learning benchmarks demonstrate the efficacy of ML-Mamba. Our model not only achieves competitive performance with other similarly sized small multimodal large-scale language models (MLLMs) but also surpasses larger MLLMs on several prominent benchmark tests, including LLaVA v1.5 [32] versions 7B and 13b.
The major contributions of this paper are three-fold:
  • We propose a novel and efficient method, i.e., ML-Mamba, which explores and utilizes multimodal learning tasks combined with the latest Mamba-2. Compared to the multimodal model adopting the original Mamba, the multimodal large-scale language model based on Mamba-2 has higher inference performance and effectiveness. Meanwhile, ML-Mamba also provides a new framework choice for multimodal large-scale language models beyond Transformer-based architectures.
  • We empirically explore the impact of different components in ML-Mamba and propose a novel multimode connector called Mamba-2 Scan Connector (MSC). MSC includes the Mamba-2 Visual Selective Scanning (MVSS) module and the SwiGLU module, which enhance representational capabilities.
  • We conduct extensive experiments on different multimodal learning benchmarks. The numerical results show that ML-Mamba achieves competitive performance compared to existing multimodal large-scale language models.
@@ 3.1 Mamba Preliminaries
By combining the modified parallel Mamba blocks with using SSD as the inner SSM layer, the Mamba-2 architecture is formed (as shown in Fig. 4(a)). The performance of Mamba-2 models of varying sizes on the Pile dataset shows that it matches or outperforms Mamba and other open-source Transformer models on standard downstream evaluations.
Figure 1: The architecture of ML-Mamba (right) uses Mamba-2 as the backbone (left). It includes a visual encoder, a multi-modal connector called the Mamba-2 Scan Connector (MSC), an MLP projector, and a language model. We use the pre-trained Mamba-2 large language model (Mamba-2 LLM) as the language model and a pre-trained visual transformer model as the visual encoder.
@@ 3.2.3 MultiModal Connector
  • MSC-MLP (Basic): It combines the multimodal connector called the Mamba-2 Scan Connector (MSC) module, which does not include the SwiGLU module and is intended to enhance the processing capability of two-dimensional non-causal visual information. Subsequently, the MLP aligns the features of vision and text (see Fig. 2(b))).
@@ 3.2.5 Training Process
We first use a 558K subset of the LAION-CC-SBU dataset to align the Mamba-2 Scan Connector (MSC) and the MLP projector. During the fine-tuning stage, we simultaneously optimized the Mamba-2 Scan Connector (MSC), the projector, and the Mamba LLM. This comprehensive training effort was executed on 8 NVIDIA A100 GPUs. The fine-tuning was conducted over two epochs, randomly sampling from the Mixed Dataset Used in LLaVA v1.5, which includes a total of 665K visual multi-round dialogue samples and pure text dialogue data.
@@ 4.1 Experimental Setup
MLP + Mamba-2 Scan Connector init Random
@@
Alignment / Fine-Tuning Samples 558K / 665K
@@
Alignment and Fine-tuning epochs 1 each
@@
ML-Mamba (ours) Mamba-2 LLM-2.7B 558K 665K 75.26 60.68 52.2 88.3 45.17 51.5
@@ 4.2 Results
To demonstrate the model’s effectiveness, we compared it with a VLM of the same scale with approximately 3B parameters, or with a larger VLM containing twice the number of parameters. As shown in Table 2, despite having only 40% of LLaVA v1.5 7B’s parameters, ML-Mamba performs comparably across multiple benchmarks and even surpasses all models on POPE.
@@ 4.3 Reasoning speed
The results from Table 3 demonstrated that although the number of image markers processed by ML-Mamba significantly increased, it still exhibited extremely fast inference speed. Compared to MobileVLM v2, although the latter has undergone multiple lightweight optimizations, the time required for ML-Mamba to complete inference is only about 30% of the former. This indicates that ML-Mamba not only maintains high speed while processing larger data, but also, thanks to the characteristics of its RNN like model, its memory usage does not significantly increase with the increase of image marker length, as such models maintain a fixed size hidden state to store historical information during the inference process.